# Spectral Persistent Homology: Persistence Signals

### 28 Dec 2023 v1

@@ 5.2 Dynamical systems (Orbit5K Dataset)
We generated 700 training and 300 testing datasets for each $\rho$ class, conducting a one-versus-one classification and using persistence diagrams for both $H_{1}$ and $H_{0}$ homologies, following the approach described in Le and Yamada (2018). We employ the same transforms with hyperparameter settings as described in the previous subsection. We adopt a regular square grid of $64\times 64$ and $128\times 128$ for all transforms in this learning task. The highest accuracies achieved with our transforms are presented in Table 2. Detailed observations for each transform across the different grids are available in Appendix B. For this dynamical particles classification task, we report the accuracy averaged over 10 runs.
In line with Carriere et al. (2020) and Hacquard (2023), we compare our methods with four kernel methods on persistence diagrams: Persistence Scale Space Kernel (PSS-K from Reininghaus et al. (2015)), Persistence Weighted Gaussian Kernel (PWG-K from Kusano et al. (2016)), Sliced Wasserstein Kernel (SW-K from Carrière et al. (2017)), and Persistence Fisher Kernel (PF-K from Le and Yamada (2018)). Moreover, we compare with two neural network architectures, Perslay and Persformer (respectively from Carriere et al. (2020) and Reinauer et al. (2021)), Persistence Images (PI from Adams et al. (2017)), and a rectangle-based classification algorithm on persistence measures (BBA from Hacquard (2023)). The accuracies in Table 2 for our competitor methods are sourced from Le and Yamada (2018) for the kernel-based approaches, and from the respective original publications for the non-kernel methods. Our methods outperform most of the kernel methods, Persistence Images, and BBA classification, but are slightly less effective than the neural networks.
${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\star}$ ${}^{\dagger}$ ${}^{\ast}$ ${}^{\star}$ ${}^{\star}$
$72.38(\pm 2.4)$ $76.63(\pm.7)$ $83.6(\pm.9)$ $85.9(\pm.8)$ $82.5$ $87.7(\pm 1.0)$ $91.2(\pm.8)$ $83.3(\pm.5)$ $84.6(\pm.8)$
@@ 6 Conclusion and further work
In this study, we introduced a new vectorization framework for persistence diagrams by treating them as signals on +$\mathbb{R}^{2}_{+}$. We employed a functional-based transforms, focusing particularly on frequency methods. This approach has proven to be stable and effective in embedding information into a vector space, preserving key topological features. Its performance, either comparable or superior to leading methods, is demonstrated across various graph classification tasks and synthetic dynamical particle classifications. The simplicity of our method, along with its potential for enhancement, makes it a strong contender in the field of TDA vectorization techniques.
However, our approach is not without limitations. A significant challenge lies in the choice of grid. Presently, the grid’s fine granularity entails high computational expense for subsequent use of our vectorization in classical ML pipelines due to high dimensionality of the embedding. Analysis of various persistence diagrams reveals a concentration of points near the origin, contrasting with extended areas of low point density. Consequently, many grid points remain unassociated with any persistence diagram points, resulting in numerous zeros in integral transforms. These zeros contribute little to no information gain, serving primarily to inflate the dimensionality of the embedding.
@@
Furthermore, our methodology is not confined to the frequency domain. Future research could explore other domains, aiming to identify additional functionals that effectively retain the information within persistence diagrams.
Another prospective direction for future work involves incorporating our vectorization concept into frameworks akin to those described in Carriere et al. (2020) or Reinauer et al. (2021). This integration could leverage neural networks or transformer architectures to refine the dimensionality of our embeddings, potentially enhancing both performance and applicability.
@@ Appendix A Proofs
$\leq\sqrt{2}\pi 4M\left\|\mathbf{x}-\mathbf{x}^{\prime}\right\|_{2}\quad\text{by Cauchy Schwarz and by $\left\|\mathbf{\omega}\right\|_{\infty}\leq\sqrt{2}M$}$
@@ Appendix B Complementary material
Dataset/Grid FOURIER Real V. GABOR WT-$db1$ WT-$db2$ WT-$db3$ WT-$coif1$ WT-$coif2$ WT-$coif3$ Mean Max Mean Max Mean Max Mean Max Mean Max Mean Max Mean Max Mean Max MUTAG 20 $87.8(\pm.8)$ $88.9$ $88.2(\pm.9)$ $89.4$ $87.8(\pm.8)$ $88.9$ $87.8(\pm.8)$ $88.9$ $88.7(\pm 1.0)$ $90.6$ $88.6(\pm.8)$ $90.0$ $87.7(\pm 1.2)$ $90.0$ $87.5(\pm.9)$ $89.4$ 32 $89.1(\pm 1.2)$ $91.1$ $89.7(\pm 1.0)$ $91.1$ $88.4(\pm 1.3)$ $90.0$ $88.6(\pm 1.0)$ $90.0$ $88.9(\pm 1.0)$ $90.6$ $88.9(\pm.7)$ $90.0$ $89.7(\pm.5)$ $90.6$ $\mathbf{89.8(\pm.9)}$ $\mathbf{91.7}$ 50 $88.1(\pm 1.1)$ $89.4$ $87.8(\pm 1.2)$ $89.4$ $86.9(\pm 1.1)$ $88.9$ $87.6(\pm.9)$ $88.9$ $88.3(\pm 1.4)$ $91.1$ $88.2(\pm.9)$ $89.4$ $88.3(\pm 1.1)$ $90.0$ $88.1(\pm.6)$ $88.9$ DHFR 20 $77.8(\pm.7)$ $78.9$ $76.8(\pm.8)$ $78.3$ $77.8(\pm.6)$ $78.9$ $78.3(\pm.7)$ $78.9$ $79.3(\pm.6)$ $80.0$ $78.5(\pm.6)$ $79.6$ $79.2(\pm 1.0)$ $81.2$ $79.0(\pm.8)$ $80.4$ 32 $79.1(\pm 1.0)$ $80.5$ $78.3(\pm.7)$ $79.5$ $81.5(\pm.7)$ $\mathbf{83.1}$ $\mathbf{81.8(\pm.7)}$ $82.9$ $81.1(\pm 1.0)$ $82.4$ $81.4(\pm.5)$ $82.1$ $80.7(\pm.5)$ $82.4$ $80.6(\pm.7)$ $81.9$ 50 $79.6(\pm.7)$ $81.2$ $78.5(\pm.8)$ $80.0$ $80.6(\pm.4)$ $81.2$ $80.2(\pm.4)$ $80.8$ $80.7(\pm.4)$ $81.5$ $81.0(\pm.5)$ $81.9$ $81.5(\pm.5)$ $82.3$ $80.9(\pm.9)$ $82.7$ COX2 20 $78.3(\pm.6)$ $78.9$ $\mathbf{80.6(\pm 1.0)}$ $\mathbf{82.4}$ $79.2(\pm.7)$ $80.4$ $79.7(\pm.7)$ $81.1$ $79.7(\pm.7)$ $80.9$ $80.0(\pm.4)$ $80.4$ $79.7(\pm.5)$ $80.4$ $79.5(\pm.8)$ $81.5$ 32 $76.7(\pm.7)$ $77.8$ $78.4(\pm.8)$ $80.0$ $77.3(\pm 1.0)$ $79.3$ $78.0(\pm 1.0)$ $79.3$ $78.0(\pm 1.1)$ $79.3$ $78.0(\pm.6)$ $78.9$ $78.2(\pm.6)$ $78.9$ $77.4(\pm.8)$ $78.5$ 50 $77.1(\pm.9)$ $78.3$ $79.5(\pm.8)$ $81.3$ $77.5(\pm.7)$ $78.5$ $78.2(\pm.5)$ $78.9$ $77.7(\pm.6)$ $78.9$ $78.7(\pm.7)$ $80.0$ $78.2(\pm 1.0)$ $79.3$ $78.8(\pm 1.2)$ $80.2$ PROTEINS 20 $72.5(\pm.6)$ $73.2$ $72.7(\pm.3)$ $73.1$ $72.4(\pm.4)$ $72.8$ $72.5(\pm.7)$ $73.4$ $72.6(\pm.6)$ $73.4$ $\mathbf{72.8(\pm.6)}$ $\mathbf{73.6}$ $72.3(\pm.5)$ $73.0$ $72.1(\pm.4)$ $72.9$ 32 $71.9(\pm.5)$ $72.8$ $72.3(\pm.5)$ $73.1$ $72.1(\pm.5)$ $72.9$ $72.3(\pm.5)$ $73.1$ $72.3(\pm.5)$ $73.3$ $72.3(\pm.6)$ $73.4$ $71.9(\pm.5)$ $73.0$ $72.1(\pm.8)$ $73.0$ 50 71.9(±.671.9($\pm$.671.9 ( ± .6 $72.7$ $72.3(\pm.6)$ $73.4$ $72.1(\pm.5)$ $73.1$ $72.4(\pm.6)$ $73.2$ $72.4(\pm.6)$ $73.2$ $72.7(\pm.5)$ $73.5$ $72.6(\pm.5)$ $73.5$ $72.6(\pm.5)$ $73.3$ IMDB-B 20 $67.6(\pm.6)$ $68.6$ $67.7(\pm.5)$ $68.6$ $67.5(\pm.8)$ $68.7$ $67.1(\pm.6)$ $68.3$ $67.0(\pm.6)$ $67.7$ $67.1(\pm.3)$ $67.5$ $66.8(\pm.7)$ $67.9$ $66.4(\pm.9)$ $68.3$ 32 $66.7(\pm.5)$ $67.7$ $67.4(\pm.7)$ $68.6$ $66.4(\pm.9)$ $68.3$ $66.9(\pm.8)$ $67.9$ $67.2(\pm.6)$ $67.9$ $66.7(\pm.8)$ $67.7$ $66.8(\pm.6)$ $67.7$ $66.8(\pm.6)$ $67.8$ 50 $67.5(\pm.8)$ $69.1$ $\mathbf{68.4(\pm.7)}$ $\mathbf{69.8}$ $68.2(\pm.6)$ $69.3$ $68.1(\pm.7)$ $69.2$ $67.4(\pm.8)$ $68.5$ $68.2(\pm.7)$ $69.1$ $67.7(\pm.4)$ $68.3$ $67.7(\pm.5)$ $68.5$ IMDB-M 20 $42.2(\pm.5)$ $43.1$ $42.0(\pm.4)$ $42.7$ $42.4(\pm.4)$ $42.9$ $42.2(\pm.5)$ $42.9$ $42.5(\pm.3)$ $43.3$ $42.3(\pm.7)$ $43.5$ $42.3(\pm.3)$ $42.8$ $42.5(\pm.4)$ $43.2$ 32 $42.8(\pm.5)$ $43.7$ $42.5(\pm.4)$ $43.2$ $42.8(\pm.5)$ $43.9$ $42.5(\pm.4)$ $43.5$ $43.1(\pm.5)$ $43.9$ $42.9(\pm.5)$ $43.6$ $42.7(\pm.4)$ $43.5$ $42.6(\pm.5)$ $43.8$ 50 $42.9(\pm.5)$ $43.7$ $42.9(\pm.2)$ $43.3$ $43.0(\pm.5)$ $43.6$ $43.3(\pm.6)$ $\mathbf{44.4}$ $43.1(\pm.5)$ $43.9$ $42.9(\pm.3)$ $43.5$ $\mathbf{43.7(\pm.5)}$ $43.9$ $43.3(\pm.4)$ $44.1$
@@
Dataset/Grid FOURIER Real V. GABOR WT-$db1$ WT-$db2$ WT-$db3$ WT-$coif1$ WT-$coif2$ WT-$coif3$ Orbit5K 64 $81.0(\pm 1.1)$ $82.2(\pm.8)$ $82.8(\pm.9)$ $80.8(\pm.7)$ $81.1(\pm.9)$ $81.0(\pm.9)$ $81.6(\pm 1.0)$ $81.3(\pm.6)$ 128 $83.4(\pm 1.0)$ $82.7(\pm.7)$ $82.2(\pm.7)$ $83.4(\pm.6)$ $83.8(\pm.7)$ $83.6(\pm 1.1)$ $83.9(\pm 1.0)$ $84.6(\pm.8)$

### 05 Jan 2024 v2

@@ 5.2 Dynamical systems (Orbit5K Dataset)
We generated 700 training and 300 testing datasets for each $\rho$ class, conducting a one-versus-one classification and using persistence diagrams for both $H_{1}$ and $H_{0}$ homologies, following the approach described in Le and Yamada (2018). We employ the same transforms with hyperparameter settings as described in the previous subsection. We adopt a regular square grid of $64\times 64$ and $128\times 128$ for all transforms in this learning task. The highest accuracies achieved with our transforms are presented in Table 2. Detailed observations for each transform across the different grids are available in Table 5. For this dynamical particles classification task, we report the accuracy averaged over 10 runs.
In line with Carriere et al. (2020) and Hacquard (2023), we compare our methods with four kernel methods on persistence diagrams: Persistence Scale Space Kernel (PSS-K from Reininghaus et al. (2015)), Persistence Weighted Gaussian Kernel (PWG-K from Kusano et al. (2016)), Sliced Wasserstein Kernel (SW-K from Carrière et al. (2017)), and Persistence Fisher Kernel (PF-K from Le and Yamada (2018)). Moreover, we compare with two neural network architectures, Perslay and Persformer (respectively from Carriere et al. (2020) and Reinauer et al. (2021)), Persistence Images (PI from Adams et al. (2017)), and a rectangle-based classification algorithm on persistence measures (BBA from Hacquard (2023)). The accuracies in Table 2 for our competitor methods are sourced from Le and Yamada (2018) for the kernel-based approaches, and from the respective original publications for the non-kernel methods. Our frequency-based methods surpass most kernel methods, Persistence Images, and BBA classification in performance, though they fall slightly behind neural networks. As for the identity transform, it outperforms some kernel methods and shows comparable results to Persistence Images. However, it generally exhibits suboptimal performance.
${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\star}$ ${}^{\dagger}$ ${}^{\ast}$ ${}^{\star}$ ${}^{\star}$ (Our) ${}^{\star}$ (Our)
$72.38(\pm 2.4)$ $76.63(\pm.7)$ $83.6(\pm.9)$ $85.9(\pm.8)$ $82.5$ $87.7(\pm 1.0)$ $91.2(\pm.8)$ $83.3(\pm.5)$ $84.6(\pm.8)$ $82.0(\pm.8)$
@@ 6 Conclusion and further work
In this study, we introduced a new vectorization framework for persistence diagrams. We employed a functional-based transforms, focusing particularly on frequency methods. This approach has proven to be stable and effective in embedding information into a vector space, preserving key topological features. Its performance, either comparable or superior to leading methods, is demonstrated across various graph classification tasks and synthetic dynamical particle classifications. The simplicity of our method, along with its potential for enhancement, makes it a strong contender in the field of TDA vectorization techniques. We also menaged to show that also with a simple identity transform, our vectorization framework is capable of reaching performance comaparable to current vectorization from the literature in TDA, questioning the direction of the field.
However, our approach is not without limitations. A significant challenge lies in the choice of grid. Presently, the grid’s fine granularity entails high computational expense for subsequent use of our vectorization in classical ML pipelines due to high dimensionality of the embedding. Due to presence of noise, persistence diagrams reveals a concentration of points near the origin, contrasting with extended areas of low point density. Consequently, many grid points remain unassociated with any persistence diagram points, resulting in numerous zeros in integral transforms. These zeros contribute little to no information gain, serving primarily to inflate the dimensionality of the embedding.
@@
Additionally, our methodology is not limited to the frequency domain. In future research, we plan to explore the use of neural network architectures for the function transform $f(\mathbf{x},\cdot)$ and learn the embedding. While this approach might reduce the direct interpretability of the embedding through qualitative analysis, it could significantly enhance both performance and applicability.
@@ Appendix A Proofs
$\leq\sqrt{2}\pi 4M\left\|\mathbf{x}-\mathbf{x}^{\prime}\right\|_{2}\quad\text{by Cauchy Schwarz and by $\left\|\|\mathbf{\omega}\|_{2}\right\|_{\infty}\leq\sqrt{2}M$}$
@@ Appendix B Complementary material
Dataset/Grid FOURIER Real V. GABOR WT-$db1$ WT-$db2$ WT-$db3$ WT-$coif1$ WT-$coif2$ WT-$coif3$ IdT Mean Max Mean Max Mean Max Mean Max Mean Max Mean Max Mean Max Mean Max Mean Max MUTAG 20 $87.8(\pm.8)$ $88.9$ $88.2(\pm.9)$ $89.4$ $87.8(\pm.8)$ $88.9$ $87.8(\pm.8)$ $88.9$ $88.7(\pm 1.0)$ $90.6$ $88.6(\pm.8)$ $90.0$ $87.7(\pm 1.2)$ $90.0$ $87.5(\pm.9)$ $89.4$ $88.1(\pm.7)$ $89.4$ 32 $89.1(\pm 1.2)$ $91.1$ $89.7(\pm 1.0)$ $91.1$ $88.4(\pm 1.3)$ $90.0$ $88.6(\pm 1.0)$ $90.0$ $88.9(\pm 1.0)$ $90.6$ $88.9(\pm.7)$ $90.0$ $89.7(\pm.5)$ $90.6$ $\mathbf{89.8(\pm.9)}$ $\mathbf{91.7}$ $89.1(\pm.9)$ $90.7$ 50 $88.1(\pm 1.1)$ $89.4$ $87.8(\pm 1.2)$ $89.4$ $86.9(\pm 1.1)$ $88.9$ $87.6(\pm.9)$ $88.9$ $88.3(\pm 1.4)$ $91.1$ $88.2(\pm.9)$ $89.4$ $88.3(\pm 1.1)$ $90.0$ $88.1(\pm.6)$ $88.9$ $87.4(\pm 1)$ $88.9$ DHFR 20 $77.8(\pm.7)$ $78.9$ $76.8(\pm.8)$ $78.3$ $77.8(\pm.6)$ $78.9$ $78.3(\pm.7)$ $78.9$ $79.3(\pm.6)$ $80.0$ $78.5(\pm.6)$ $79.6$ $79.2(\pm 1.0)$ $81.2$ $79.0(\pm.8)$ $80.4$ $78.7(\pm.9)$ $79.7$ 32 $79.1(\pm 1.0)$ $80.5$ $78.3(\pm.7)$ $79.5$ $81.5(\pm.7)$ $\mathbf{83.1}$ $\mathbf{81.8(\pm.7)}$ $82.9$ $81.1(\pm 1.0)$ $82.4$ $81.4(\pm.5)$ $82.1$ $80.7(\pm.5)$ $82.4$ $80.6(\pm.7)$ $81.9$ $81.1(\pm.5)$ $82.0$ 50 $79.6(\pm.7)$ $81.2$ $78.5(\pm.8)$ $80.0$ $80.6(\pm.4)$ $81.2$ $80.2(\pm.4)$ $80.8$ $80.7(\pm.4)$ $81.5$ $81.0(\pm.5)$ $81.9$ $81.5(\pm.5)$ $82.3$ $80.9(\pm.9)$ $82.7$ $81.2(\pm.7)$ $82.7$ COX2 20 $78.3(\pm.6)$ $78.9$ $\mathbf{80.6(\pm 1.0)}$ $\mathbf{82.4}$ $79.2(\pm.7)$ $80.4$ $79.7(\pm.7)$ $81.1$ $79.7(\pm.7)$ $80.9$ $80.0(\pm.4)$ $80.4$ $79.7(\pm.5)$ $80.4$ $79.5(\pm.8)$ $81.5$ $79.4(\pm.7)$ $80.4$ 32 $76.7(\pm.7)$ $77.8$ $78.4(\pm.8)$ $80.0$ $77.3(\pm 1.0)$ $79.3$ $78.0(\pm 1.0)$ $79.3$ $78.0(\pm 1.1)$ $79.3$ $78.0(\pm.6)$ $78.9$ $78.2(\pm.6)$ $78.9$ $77.4(\pm.8)$ $78.5$ $77.5(\pm.5)$ $78.5$ 50 $77.1(\pm.9)$ $78.3$ $79.5(\pm.8)$ $81.3$ $77.5(\pm.7)$ $78.5$ $78.2(\pm.5)$ $78.9$ $77.7(\pm.6)$ $78.9$ $78.7(\pm.7)$ $80.0$ $78.2(\pm 1.0)$ $79.3$ $78.8(\pm 1.2)$ $80.2$ $78.0(\pm.4)$ $78.5$ PROTEINS 20 $72.5(\pm.6)$ $73.2$ $72.7(\pm.3)$ $73.1$ $72.4(\pm.4)$ $72.8$ $72.5(\pm.7)$ $73.4$ $72.6(\pm.6)$ $73.4$ $\mathbf{72.8(\pm.6)}$ $\mathbf{73.6}$ $72.3(\pm.5)$ $73.0$ $72.1(\pm.4)$ $72.9$ $71.9(\pm.7)$ $73.2$ 32 $71.9(\pm.5)$ $72.8$ $72.3(\pm.5)$ $73.1$ $72.1(\pm.5)$ $72.9$ $72.3(\pm.5)$ $73.1$ $72.3(\pm.5)$ $73.3$ $72.3(\pm.6)$ $73.4$ $71.9(\pm.5)$ $73.0$ $72.1(\pm.8)$ $73.0$ $72.2(\pm.5)$ $73.2$ 50 71.9(±.671.9($\pm$.671.9 ( ± .6 $72.7$ $72.3(\pm.6)$ $73.4$ $72.1(\pm.5)$ $73.1$ $72.4(\pm.6)$ $73.2$ $72.4(\pm.6)$ $73.2$ $72.7(\pm.5)$ $73.5$ $72.6(\pm.5)$ $73.5$ $72.6(\pm.5)$ $73.3$ $71.4(\pm.4)$ $72.1$ IMDB-B 20 $67.6(\pm.6)$ $68.6$ $67.7(\pm.5)$ $68.6$ $67.5(\pm.8)$ $68.7$ $67.1(\pm.6)$ $68.3$ $67.0(\pm.6)$ $67.7$ $67.1(\pm.3)$ $67.5$ $66.8(\pm.7)$ $67.9$ $66.4(\pm.9)$ $68.3$ $67.2(\pm.7)$ $68.3$ 32 $66.7(\pm.5)$ $67.7$ $67.4(\pm.7)$ $68.6$ $66.4(\pm.9)$ $68.3$ $66.9(\pm.8)$ $67.9$ $67.2(\pm.6)$ $67.9$ $66.7(\pm.8)$ $67.7$ $66.8(\pm.6)$ $67.7$ $66.8(\pm.6)$ $67.8$ $66.5(\pm.6)$ $67.9$ 50 $67.5(\pm.8)$ $69.1$ $\mathbf{68.4(\pm.7)}$ $\mathbf{69.8}$ $68.2(\pm.6)$ $69.3$ $68.1(\pm.7)$ $69.2$ $67.4(\pm.8)$ $68.5$ $68.2(\pm.7)$ $69.1$ $67.7(\pm.4)$ $68.3$ $67.7(\pm.5)$ $68.5$ $66.7(\pm.9)$ $67.9$ IMDB-M 20 $42.2(\pm.5)$ $43.1$ $42.0(\pm.4)$ $42.7$ $42.4(\pm.4)$ $42.9$ $42.2(\pm.5)$ $42.9$ $42.5(\pm.3)$ $43.3$ $42.3(\pm.7)$ $43.5$ $42.3(\pm.3)$ $42.8$ $42.5(\pm.4)$ $43.2$ $42.8(\pm.4)$ $43.6$ 32 $42.8(\pm.5)$ $43.7$ $42.5(\pm.4)$ $43.2$ $42.8(\pm.5)$ $43.9$ $42.5(\pm.4)$ $43.5$ $43.1(\pm.5)$ $43.9$ $42.9(\pm.5)$ $43.6$ $42.7(\pm.4)$ $43.5$ $42.6(\pm.5)$ $43.8$ $43.1(\pm.5)$ $44.3$ 50 $42.9(\pm.5)$ $43.7$ $42.9(\pm.2)$ $43.3$ $43.0(\pm.5)$ $43.6$ $43.3(\pm.6)$ $\mathbf{44.4}$ $43.1(\pm.5)$ $43.9$ $42.9(\pm.3)$ $43.5$ $\mathbf{43.7(\pm.5)}$ $43.9$ $43.3(\pm.4)$ $44.1$ $43.1(\pm.4)$ $43.7$
@@
Dataset ${}^{\dagger}$ ${}^{\star}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\star}$ ${}^{\star}$ ${}^{\dagger}$ LITE (Our) LITE-IdT (Our)
${}^{\star}$ ${}^{\dagger}$ ${}^{\star}$ ${}^{\dagger}$
MUTAG $88.3$ $90.3$ $92.1$ $86.7$ $89.0$ $79.2$ $86.1$ $89.8$ $88.3$ $90.4$ $89.8$ $91.7$ $89.2$ $90.7$
COX2 $78.4$ $81.4$ $76.0$ $79.9$ $80.9$ $79.4$ $81.2$ $80.6$ $82.4$ $79.4$ $80.4$
PROTEINS $72.6$ $75.8$ $73.4$ $76.3$ $75.8$ $65.4$ $67.5$ $74.8$ $71.4$ $74.7$ $72.8$ $73.6$ $72.2$ $73.2$
DHFR $78.4$ $81.5$ $70.9$ $81.7$ $80.3$ $82.7$ $80.5$ $81.8$ $83.1$ $81.2$ $82.7$
IMDB-B $72.9$ $71.9$ $73.6$ $73.1$ $74.3$ $54.0$ $68.7$ $71.2$ $74.8$ $69.4$ $68.4$ $69.8$ $67.2$ $68.3$
IMDB-M $50.3$ $47.7$ $52.4$ $50.3$ $52.1$ $36.3$ $46.9$ $48.8$ $47.8$ $46.7$ $43.7$ $44.4$ $43.1$ $44.3$
Dataset/Grid FOURIER Real V. GABOR WT-$db1$ WT-$db2$ WT-$db3$ WT-$coif1$ WT-$coif2$ WT-$coif3$ IdT Orbit5K 64 $81.0(\pm 1.1)$ $82.2(\pm.8)$ $\mathbf{82.8(\pm.9)}$ $80.8(\pm.7)$ $81.1(\pm.9)$ $81.0(\pm.9)$ $81.6(\pm 1.0)$ $81.3(\pm.6)$ $82.0(\pm.8)$ 128 $83.4(\pm 1.0)$ $82.7(\pm.7)$ $82.2(\pm.7)$ $83.4(\pm.6)$ $83.8(\pm.7)$ $83.6(\pm 1.1)$ $83.9(\pm 1.0)$ $\mathbf{84.6(\pm.8)}$ $78.6(\pm 1.1)$